### v2

# sources.bib

Abstract.
Let $F$ be a tensor of degree $d$ over a field $K.$ Write $\textnormal{prk}(F)$ for the partition rank of $F$ and $\overline{\textnormal{prk}}(F):=\textnormal{prk}(F\otimes_{K}\overline{K}),$ where $\overline{K}$ is an algebraic closure of $K.$ We prove that $\textnormal{prk}(F)\ll_{d,K}\overline{\textnormal{prk}}(F),$ as long as $K$ is both infinite and of finite absolute transcendence degree, partially resolving a conjecture of Adiprasito, Kazhdan and Ziegler. For finite fields we prove a result which, together with results of Moshkovitz-Zhu, implies $\textnormal{prk}(F)\ll_{d}\overline{\textnormal{prk}}(F)\log(\overline{\textnormal{prk}}(F)+1).$ This improves the state of the art bound for partition rank in terms of analytic rank. By a standard argument, we deduce similar bounds for strength of degree $d$ polynomials, as long as $\textnormal{char}(K)>d.$
BB was supported by NSF grant DMS-2101075
AL was supported by NSF grant DMS-2402041
1. Introduction
In $1985,$ Schmidt [Sch] introduced the notion of strength for polynomials, proceeded to derive some remarkable arithmetic and geometric properties of polynomials with high strength, and then used these properties to obtain counting results for integer solutions to systems of homogeneous polynomial equations. In the following decades, strength has found myriad applications in such diverse fields as additive combinatorics, analytic number theory and commutative algebra. It has evolved into a central tool for understanding the behavior of polynomials of fixed degree in many variables. In 2009, motivated by the inverse conjecture for the Gowers norm, Green and Tao [GT] proved that polynomials of high strength over finite fields have equidistributed values. Their result was extended and improved by many others, see e.g. [HS], [TZ], [KL], [BL], [Mil], [Jan], [CM] and [MZ]. In 2020, Ananyan and Hochster [AH] proved Stillman’s conjecture regarding the projective dimension of ideals in polynomial rings. Strength played a crucial role in their proof, as well as in subsequent results, such as [ESS1], [Dra], [BDE] and [ESS2].
In the special case of tensors, strength is equivalent to another invariant, partition rank, introduced by Naslund [Nas]. For algebraically closed fields, partition rank in turn is closely related to a quantity called geometric rank, introduced in [KMZ20]. Geometric rank and its close cousin Birch rank play a key role in various applications of the Hardy-Littlewood circle method: see e.g. [Bir], [CMag], [BHB] and [Ski]. Since partition rank is an arithmetic quantity, it can potentially decrease under field extensions. In arithmetic applications of strength, such as [BDS-Brauer] and [LS], it’s important to bound the possible decrease in strength independently of the number of variables. See [LZ] for several further applications of such a bound.
A bound of this type was obtained for cubics in [Der] and [AKZ], for quartics in [KP] and for arbitrary degree in [LZ], albeit only for some fields. More general results were obtained in [BDLZ] and [BDS]. The strongest conjecture in this direction, conjecture 1.1, was made by Adiprasito, Kazhdan and Ziegler in [AKZ]. Thus far, this has remained wide open for degrees greater than 3. Theorem 1.3 constitutes a partial resolution of this conjecture. Before stating our results, we introduce some notation.
Let $K$ be a field with algebraic closure $\overline{K}$, let $V$ be a finite-dimensional vector space over $K$ and let $F:V^{d}\to K$ be a multi-linear map, i.e. a tensor of degree $d.$ Write $\textnormal{prk}(F)$ for the partition rank of $F$ over $K$ (defined in §2) and $\overline{\textnormal{prk}}(F)$ for the partition rank of $F$ over $\overline{K}.$ The inequality $\overline{\textnormal{prk}}(F)\leq\textnormal{prk}(F)$ is immediate from the definition. In [AKZ], it was conjectured that the reverse inequality also holds, up to a constant factor depending only on the degree.
Conjecture 1.1.
$\textnormal{prk}(F)\ll_{d}\overline{\textnormal{prk}}(F).$
Our progress on this conjecture, theorem 1.3, follows from a result of independent interest which relates other kinds of rank for tensors. Write $\textnormal{grk}(F)$ for the geometric rank of $F$ and $\textnormal{ark}(F)$ for the analytic/arithmetic rank of $F$, when $K$ is finite/infinite, respectively. All of these ranks will be defined in §$\S 2$.
Remark.
We use $\log$ throughout the paper to denote the base-$2$ logarithm.
Theorem 1.2.
  1. (1)
    For infinite $K$ we have
    $\textnormal{grk}(F)\leq\textnormal{ark}(F)\leq(d-1)^{r}\cdot\textnormal{grk}(F),$
    where
    $r(K)=\begin{cases}\textnormal{trdeg}(K/\mathbf{Q})+1\ ,\ &\textnormal{char}(K)=0,\\ \max(\textnormal{trdeg}(K/\mathbf{F}_{p}),1),\ &\textnormal{char}(K)=p>0.\end{cases}$
  2. (2)
    For finite $K$ we have
    $\frac{\textnormal{grk}(F)}{2\cdot(2\log d+1)^{d-1}}\leq\textnormal{ark}(F)\leq(d-1)\cdot\textnormal{grk}(F).$
Together with the results of [KLP] and [MZ], theorem 1.2 implies the following.
Theorem 1.3.
  1. (1)
    For infinite $K$ we have
    $\textnormal{prk}(F)\leq(2^{d-1}-1)(d-1)^{r}\overline{\textnormal{prk}}(F),$
    where $r=r(K)$ is the constant of theorem 1.2.
  2. (2)
    For finite $K$ we have
    $\textnormal{prk}(F)\ll_{d}\overline{\textnormal{prk}}(F)\cdot\log(\overline{\textnormal{prk}}(F)+1).$
The related question of bounds for partition rank in terms of analytic rank began with the work of Green and Tao [GT]. Notable milestones include polynomial bounds obtained independently by Milićević [Mil] and Janzer [Jan], linear bounds for large fields due to Cohen-Moshkovitz [CM] and the current record of quasi-linear bounds due to Moshkovitz-Zhu [MZ]. Our theorem 1.3 combines with the results of Moshkovitz-Zhu to improves their bound further.
Corollary 1.4.
For finite $K$ we have
$\textnormal{prk}(F)\ll_{d}\textnormal{ark}(F)\cdot\log(\textnormal{ark}(F)+1).$
By standard arguments, theorem 1.3 implies a similar inequality for strength of polynomials in high characteristic. For a form $f\in K[V]$ of degree $d,$ write $\textnormal{str}(f)$ for the strength of $f$ over $K$ and $\overline{\textnormal{str}}(f)$ for the strength of $f$ over $\overline{K},$ where strength is defined in §$\S 2.$
Corollary 1.5.
Suppose that $\textnormal{char}(K)=0$ or $\textnormal{char}(K)>d.$
  1. (1)
    If $K$ is infinite then
    $\textnormal{str}(f)\leq(2^{d-1}-1)(d-1)^{r}\binom{d}{\lfloor d/2\rfloor}\cdot\overline{\textnormal{str}}(f),$
    where $r=r(K)$ is defined as in theorem 1.3.
  2. (2)
    If $K$ is finite then
    $\textnormal{str}(f)\ll_{d}\overline{\textnormal{str}}(f)\log(\overline{\textnormal{str}}(f)+1).$
Remark.
For finite fields the above results were recently proved in [ChenYe], [MZ-grar] and [MZ] via a different method. The bounds obtained there are slightly weaker.
Structure of the proof
The proof of theorem 1.2 begins by applying Weil restriction to reduce to the case of purely transcendental extensions of either $\mathbf{Q}$ or $\mathbf{F}_{q}.$ By choosing a basis for $V$ we may assume without loss of generality that $F$ has coefficients in $A,$ where $A=\mathbf{Z}[t_{1},\ldots,t_{s}]$ or $A=\mathbf{F}_{q}[t_{1},\ldots,t_{s}],$ respectively. The quotients of $A$ by maximal ideals are then finite fields, and $F$ reduces to a multi-linear form over each of these finite fields. We proceed to characterize geometric rank via the asymptotic count of singular points for these reductions. By a scaling inequality, we deduce that many of these singular points have small height and therefore lift to genuine singular points of $F$. For finite fields, there is one final step consisting of estimating the size of the fibers of the ”evaluation at zero” map, which sends singular points with coefficients in $\mathbf{F}_{q}[t]$ to ones with coefficients in $\mathbf{F}_{q}.$
Layout of the paper
In §$\S 2$ we give the definitions of the various notions of rank in consideration and prove that theorem 1.2 implies theorem 1.3 and corollaries 1.4 and 1.5. We also recall the proof of the inequalities in theorem 1.2 which are already known. In §$\S 3$ we reduce the proof of theorem 1.2 to purely transcendental extensions of $\mathbf{Q}$ or $\mathbf{F}_{q}$ and establish results relating point counting to dimension, including an asymptotic characterization of geometric rank. In §$\S 4$ we complete the proof of theorem 1.2 in positive characteristic and in §$\S 5$ we prove it in characteristic zero.
Acknowledgements
The first author is grateful to Karen Smith for feedback on §3. The second author is grateful to Andrew Snowden for suggesting the Weil restriction argument and to Guy Moshkovitz for enlightening conversations.
2. Preliminaries
2.1. Definitions
As before, let $V$ be a finite-dimensional vector space over $K$ and $F:V^{d}\to K$ a multi-linear map. For any field extension $L/K,$ write $F_{L}:(V\otimes L)^{d}\to L$ for the multi-linear form corresponding to $F.$ We begin by recalling the definition of partition rank, due to Naslund [Nas].
Definition 2.1 (Partition rank).
We say that $F$ has partition rank one if there exists a subset $\varnothing\neq I\subsetneq[d]$ such that $F(x)=G((x_{i})_{i\in I})\cdot H((x_{j})_{j\in[d]\setminus I})$, where $G:V^{|I|}\to K$ and $H:V^{d-|I|}\to K$ are nonzero multi-linear forms. In general, the partition rank of $F$ is
$\textnormal{prk}(F)=\min\{r:F=F_{1}+\ldots+F_{r}\},$
where the $F_{i}:V^{d}\to K$ are multi-linear forms of partition rank one. Write $\overline{\textnormal{prk}}(F)$ for $\textnormal{prk}(F_{\overline{K}}),$ where $\overline{K}$ is an algebraic closure of $K.$
Geometric rank was introduced by Kopparty, Moshkovitz and Zuiddam [KMZ20].
Definition 2.2 (Geometric rank).
The geometric rank of $F$ is $\textnormal{grk}(F)=\operatorname{codim}_{V^{d-1}}S_{F},$ where
$S_{F}=(x\in V^{d-1}:F(x,\cdot)\equiv 0).$
Remark.
For fixed $r,$ the tensors of geometric rank $\leq r$ form a Zariski closed subset [KMZ20]. Together with lemma 2.5, this means that geometric rank is closely related to border rank.
For finite fields, analytic rank is an arithmetic analogue of geometric rank, introduced by Gowers and Wolf in [GWol]. We extend the analogy to infinite fields via a quantity which we call arithmetic rank.
Definition 2.3 (Analytic and arithmetic rank).
  1. (1)
    For $K=\mathbf{F}_{q},$ the analytic rank of $F$ is
    $\textnormal{ark}(F)=-\log_{q}\frac{|S_{F}(K)|}{q^{(d-1)\dim V}}.$
  2. (2)
    For infinite $K,$ the arithmetic rank is
    $\textnormal{ark}(F)=\operatorname{codim}_{V^{d-1}}\overline{S_{F}(K)},$
    where the closure is taken with respect to the Zariski topology.
Lemma 2.4.
Let $K$ be an infinite field and $L/K$ a field extension. Let $F$ be a tensor defined over $K$. Then $\textnormal{ark}(F)\geq\textnormal{ark}(F_{L}).$
Proof.
We have
$\overline{S_{F}(K)}\times_{\operatorname{Spec}K}\operatorname{Spec}L\subseteq\overline{(S_{F}\times_{\operatorname{Spec}K}\operatorname{Spec}L)(L)}$
Taking codimensions of both sides, we conclude the result. ∎
2.2. Proof of theorem 1.3 and corollary 1.4
In [KLP], Kazhdan, Polishchuk and the second author proved the following result.
Lemma 2.5.
For any infinite field $K,$
$\textnormal{grk}(F)\leq\overline{\textnormal{prk}}(F)\leq\textnormal{prk}(F)\leq(2^{d-1}-1)\textnormal{ark}(F).$
Proof of theorem 1.3 for infinite fields.
Combining this lemma with theorem 1.2 yields
$\textnormal{prk}(F)$ $\leq(2^{d-1}-1)\textnormal{ark}(F)\leq(2^{d-1}-1)(d-1)^{r}\cdot\textnormal{grk}(F)$
$\leq(2^{d-1}-1)(d-1)^{r}\cdot\overline{\textnormal{prk}}(F).$
Now, suppose that $K=\mathbf{F}_{q}$ is a finite field. We abbreviate $F_{l}=F_{\mathbf{F}_{q^{l}}}$ and note the following simple inequality, proved in [CM, Proof of corollary 1].
Lemma 2.6.
We have $\textnormal{prk}(F)\leq l\cdot\textnormal{prk}(F_{l}).$
The heavy lifting will be done by the following result of Moshkovitz and Zhu [MZ].
Proposition 2.7.
There exists a constant $C=C(d)$ such that the following holds. For any $a\geq 1,$ if $\textnormal{ark}(F_{l})\leq a$ and $q^{l}\geq Ca^{d-2}$ then $\textnormal{prk}(F_{l})\leq 2^{d-1}a.$
Proof of theorem 1.3 and corollary 1.4 for finite fields.
Writing $r=\overline{\textnormal{prk}}(F),$ we have $\textnormal{grk}(F)\leq r$ by lemma 2.5 and so $\textnormal{ark}(F_{l})\leq(d-1)r$ for all $l$ by theorem 1.2. Let
$l=\lceil\log C+(d-2)\left(\log(d-1)+\log r\right)\rceil,$
where $C=C(d)$ is the constant of proposition 2.7. Note that $l\ll_{d}\log(r+1).$ Our choice of $l$ guarantees that $q^{l}\geq 2^{l}\geq C[(d-1)r]^{d-2}$ so we may apply proposition 2.7 with $a=(d-1)r$ to obtain
$\textnormal{prk}(F_{l})\leq 2^{d-1}(d-1)r.$
Applying lemma 2.6 gives
$\textnormal{prk}(F)\leq l\cdot 2^{d-1}(d-1)r\ll_{d}r\log(r+1)=\overline{\textnormal{prk}}(F)\cdot\log(\overline{\textnormal{prk}}(F)+1).$
To prove corollary 1.4, combine theorems 1.2 and 1.3 with lemma 2.5 to get
$\textnormal{prk}(F)$ $\ll_{d}\overline{\textnormal{prk}}(F)\cdot\log(\overline{\textnormal{prk}}(F)+1)\ll_{d}\textnormal{grk}(F)\cdot\log(\textnormal{grk}(F)+1)$
$\ll_{d}\textnormal{ark}(F)\cdot\log(\textnormal{ark}(F)+1).$
2.3. Polarization and proof of corollary 1.5
Let $f\in K[V]$ be a form of degree $d.$
Definition 2.8 (Strength).
The strength of $f$ is
$\textnormal{str}(f)=\min\left\{s:f=\sum_{i=1}^{s}g_{i}h_{i}\right\},$
where $g_{i},h_{i}\in K[V]$ are forms of degree less than $d$. Write $\overline{\textnormal{str}}(f)$ for the strength of $f$ when considered as a form over $\overline{K}$.
We assume now that $\textnormal{char}(K)=0$ or $\textnormal{char}(K)>d.$ There is an associated symmetric multi-linear form $F:V^{d}\to K$ given by
$F(h_{1},\ldots,h_{d})=\nabla_{h_{1}}\ldots\nabla_{h_{d}}f(0).$
It is easy to see that the strength of $f$ and the partition rank of $F$ are closely related, see e.g. [LZ, Claim 3.2].
Lemma 2.9.
We have
$\textnormal{str}(f)\leq\textnormal{prk}(F)\leq\binom{d}{\lfloor d/2\rfloor}\textnormal{str}(f),$
and the same holds for $\overline{\textnormal{str}}(f)$ vis-à-vis $\overline{\textnormal{prk}}(F)$.
Proof of corollary 1.5.
Apply theorem 1.3 and lemma 2.9. For infinite fields, we get
$\textnormal{str}(f)$ $\leq\textnormal{prk}(F)\leq(2^{d-1}-1)(d-1)^{r}\cdot\overline{\textnormal{prk}}(F)$
$\leq(2^{d-1}-1)(d-1)^{r}\binom{d}{\lfloor d/2\rfloor}\cdot\overline{\textnormal{str}}(f).$
For finite fields, we get
$\textnormal{str}(f)$ $\leq\textnormal{prk}(F)\ll_{d}\overline{\textnormal{prk}}(F)\cdot\log(\overline{\textnormal{prk}}(F)+1)$
$\ll_{d}\overline{\textnormal{str}}(F)\cdot\log(\overline{\textnormal{str}}(F)+1).$
2.4. Proof of known inequalities in theorem 1.2
For infinite fields, $\textnormal{grk}(F)\leq\textnormal{ark}(F)$ is an immediate consequence of the definitions. The inequality $\textnormal{grk}(F)\ll_{d}\textnormal{ark}(F)$ for finite fields was essentially established in [AKZ]; we recall the proof here for the reader’s convenience.
We begin with the following result from [AKZ]:
Lemma 2.10.
If $K=\mathbf{F}_{q}$ and $q\geq d^{2}$ Then
$\textnormal{grk}(F)\leq 2\cdot\textnormal{ark}(F).$
The restriction on the field size can be handled via a result of Moshkovitz-Cohen [CM].
Lemma 2.11.
We have $\textnormal{ark}(F_{l})\leq l^{d-1}\textnormal{ark}(F).$
Choosing $l=\lceil 2\log d\rceil,$ we have $q^{l}\geq 2^{l}\geq d^{2}$ so we can apply lemmas 2.10 and 2.11 to obtain
$\textnormal{grk}(F)$ $=\textnormal{grk}(F_{l})\leq 2\cdot\textnormal{ark}(F_{l})$
$\leq 2l^{d-1}\textnormal{ark}(F)\leq 2\cdot(2\log d+1)^{d-1}\textnormal{ark}(F),$
as desired.
3. Weil restriction, dimension and Point Counting
Definition 3.1.
Let $K$ be a field. Let $D_{K,d}$ denote the least positive integer such that for any finite dimensional vector space $V$ over $K$ and any multi-linear map $G:V^{d-1}\to V$, we have
(1) $\operatorname{codim}_{V^{d-1}}\overline{Z_{G}(K)}\leq D_{K,d}\cdot\operatorname{codim}_{V^{d-1}}Z_{G},$
where $Z_{G}=(G(x)=0).$
The point of this section is to show how we can relate the constants $D_{K,d},D_{L,d}$ when $L$ is an extension field or a residue field of $K$.
3.1. Weil Restriction
Our first result deals with finite separable extensions.
Lemma 3.2 (Weil restriction).
Let $L/K$ be a finite separable extension of infinite fields. Then $D_{L,d}\leq D_{K,d}.$
Proof.
Let $\ell=[L:K].$ Let $V$ be a finite-dimensional vector space over $L$ and $G:V^{d-1}\to V$ a multi-linear form. Denote by $V_{K}$ the $K$-vector space (of dimension $\ell\cdot\dim V$) and by $G_{K}:V_{K}^{d-1}\to V_{K}$ the multi-linear form (over $K$) obtained by restricting scalars. By [BLR90], §7.6, Theorem 4, we have:
(2) $R_{L/K}(Z_{G})\cong Z_{G_{K}}.$
By [JLMM22], Corollary 3.8 we have $\ell\cdot\operatorname{codim}_{V^{d-1}}(Z_{G})=\operatorname{codim}_{V_{K}^{d-1}}(Z_{G_{K}})$. By the definitions of Zariski closure and Weil restriction, we have $Z_{G_{K}}(K)\subset R_{L/K}(\overline{Z_{G}(L)}).$ Moreover, $R_{L/K}(\overline{Z_{G}(L)})$ is closed in $V_{K}^{d-1}$ ([BLR90], §7.6, Proposition 2), so
$\overline{Z_{G_{K}}(K)}\subset R_{L/K}(\overline{Z_{G}(L)}).$
Taking codimensions, we obtain the inequality
(3) $\operatorname{codim}_{V_{K}^{d-1}}\overline{Z_{G_{K}}(K)}\geq\ell\cdot\operatorname{codim}_{V^{d-1}}\overline{Z_{G}(L)}.$
We may now apply Equations 1, 2 and 3 to obtain:
$\operatorname{codim}\overline{Z_{G}(L)}\leq\frac{\operatorname{codim}\overline{Z_{G_{K}}(K)}}{\ell}\leq\frac{D_{K,d}\operatorname{codim}Z_{G_{K}}}{\ell}=D_{K,d}\operatorname{codim}Z_{G}.$
Next, we may reduce the proof of Theorem 1.2 (1) to the case of purely transcendental extensions of $\mathbf{Q}$ or $\mathbf{F}_{q}$.
Proposition 3.3 (Reduction to Pure Transcendental Case).
Suppose Theorem 1.2 (1) holds for the fields $K=\mathbf{Q},\ \mathbf{F}_{q}(t)$ and their purely transcendental extensions, and suppose Theorem 1.2 (2) holds. Then Theorem 1.2 (1) holds.
Proof.
We consider three types of infinite field $K$:
  • (i)
    A field of infinite transcendence degree over $\mathbf{Q}$ or $\mathbf{F}_{q}$
  • (ii)
    A field of finite transcendence degree over $\mathbf{Q}$ or $\mathbf{F}_{q}(t)$
  • (iii)
    An infinite field contained in $\overline{\mathbf{F}_{p}}$ for some $p$.
In case (i), the result is trivial (and also meaningless).
For case (ii), let $F$ be a $d$-tensor with coefficients in $K$. Let $K_{0}\subseteq K$ denote the subfield generated by the coefficients of $F$ over the prime field of $K$. By construction, the parameter $r$ satisfies $r(K_{0})\leq r(K).$ Let $V_{0}$ be a $K_{0}$-vector space and $F_{0}:(V_{0})^{d}\to K_{0}$ such that $(F_{0})_{K}\cong F$.
Every finitely-generated extension of a perfect field admits a separating transcendence basis (see [AAA], Theorem 3 or [vdW], Lemma 1), so in particular, $K_{0}$ is isomorphic to a finite separable extension of the rational function field $\mathbf{Q}(t_{1},\dots,t_{r(K_{0})-1})$ or $\mathbf{F}_{q}(t_{1},\dots,t_{r(K_{0})})$. By assumption, Theorem 1.2 holds for $\mathbf{Q}(t_{1},\dots,t_{r(K_{0})-1})$ and $\mathbf{F}_{q}(t_{1},\dots,t_{r(K_{0})})$, so Lemma 3.2 implies that Theorem 1.2 holds for $K_{0}$. Consequently, we have
$\textnormal{ark}(F)\leq\textnormal{ark}(F_{0})\leq(d-1)^{r(K_{0})}\textnormal{grk}(F_{0})\leq(d-1)^{r(K)}\textnormal{grk}(F_{0})=(d-1)^{r(K)}\textnormal{grk}(F).$
We will return to the proof of case (iii) after a brief interlude with necessary results. ∎
Definition 3.4 (Complexity).
Suppose an affine variety $Z\subseteq\mathbb{A}^{N}$ is cut out (set theoretically) by polynomials $P_{1},\dots,P_{m}$. Then the complexity of $Z$ is equal to $\max(N,m,\deg(P_{1}),\dots,\deg(P_{m}))$.
Lemma 3.5 ([T12], Lemma 1 or [LW], Lemma 1).
Let $Z$ be an affine variety of complexity at most $M$ defined over $\mathbf{F}_{q}$. Then
$|Z(\mathbf{F}_{q})|\ll_{M}q^{\dim Z}.$
Remark.
Lang and Weil ([LW]) gave implied constants for the bounds in Lemma 3.5 and Lemma 3.8) in terms of three parameters: $\dim Z,\deg Z,N$. As in [T12], we combine these three parameters into one parameter: complexity. While this new parameter is apparently less robust, since we do not care about the implied constants from these two results, Definition 3.4 allows us to simplify the rest of our arguments in this section.
As a consequence of Lemma 3.5, we are able to prove case (iii) of 3.3.
End of proof of 3.3.
Fix a multi-linear form $F$ over $K\subseteq\overline{\mathbf{F}_{p}}$. The coefficients of $F$ lie in some finite subfield $\mathbf{F}_{q}$ for some $q=p^{e}$. Let $V^{\prime}$ be an $\mathbf{F}_{q}$-vector space and let $F^{\prime}:(V^{\prime})^{d}\to\mathbf{F}_{q}$ be a multi-linear form such that $F^{\prime}_{K}\cong F$. Write $K=\bigcup_{i>0}\mathbf{F}_{q^{e_{i}}}$, where $1=e_{1}<e_{2}<\dots$ is an increasing sequence of positive integers. Applying Lemma 3.5 to $Z=\overline{S_{F}(K)}$, we obtain
$\dim Z\geq\lim\sup_{i\to\infty}\log_{q^{e_{i}}}|Z(\mathbf{F}_{q^{e_{i}}})|,$
or equivalently
(4) $\textnormal{ark}(F)\leq\lim\inf_{i\to\infty}\textnormal{ark}(F^{\prime}_{e_{i}}).$
By Theorem 1.2 (2), the right-hand side of Equation 4 is bounded above by $(d-1)\textnormal{grk}(F)$, proving the claim. ∎
3.2. Reduction to Finite Fields and the Lang-Weil Bound
Kopparty, Moshkovitz, and Zuiddam proved the following result which connects the geometric rank of a $d$-tensor over $\mathbf{Q}$ to the analytic ranks of a family of prime characteristic models.
Proposition 3.6 ([KMZ20], Theorem 8.1).
Let $F:(\mathbf{Z}^{n})^{d}\to\mathbf{Z}$ be a $d$-tensor. Then we have
$\liminf_{p\to\infty}\textnormal{ark}(F\otimes_{\mathbf{Z}}\mathbf{F}_{p})=\textnormal{grk}(F\otimes_{\mathbf{Z}}\mathbf{Q}).$
For our purposes, we require an analogous result for a $d$-tensor $F:(\mathbf{F}_{q}[t]^{n})^{d}\to\mathbf{F}_{q}[t]$. The remainder of this section will be devoted to the proof of the following proposition.
Proposition 3.7.
Let $A=\mathbf{F}_{q}[t],K=\mathbf{F}_{q}(t)$ and let $F:(A^{n})^{d}\to A$ be a $d$-tensor. We have
(5) $\liminf_{\mathfrak{p}\in\max\operatorname{Spec}A,|A/\mathfrak{p}|\to\infty}\textnormal{ark}(F\otimes_{A}A/\mathfrak{p})=\textnormal{grk}(F\otimes_{A}K).$
Lemma 3.8 ([T12], Theorem 2 or [LW], Theorem 1).
Suppose $Z$ is an affine variety of complexity at most $M$ defined over $\mathbf{F}_{q}.$ Let $c=c(Z)$ be the number of top-dimensional geometrically irreducible components of $Z.$ Then
$|Z(\mathbf{F}_{q})|=(c+O_{M}(q^{-1/2}))q^{\dim Z}.$
We are now ready to begin the setup for 3.7. Let $A=\mathbf{F}_{q}[t]$ for a finite field $\mathbf{F}_{q}$. Let $K=\mathbf{F}_{q}(t)=\text{Frac}(A)$. A variety $X/K$ can be viewed as the generic point of a finite type $A$-scheme $\mathfrak{X}$. In this subsection, we’ll build up results that will allow us to connect properties of $X$ to properties of closed fibers of $\mathfrak{X}$ over $\operatorname{Spec}(A)$.
The next two propositions are standard results on constructible properties of schemes.
Lemma 3.9 ([stacks-project], Tag 0559).
Let $X\to Y$ be a morphism of schemes. Assume
  • $Y$ is irreducible with generic point $\eta$,
  • $X_{\eta}$ is geometrically irreducible,
  • $X\to Y$ is finite type.
Then there exists a nonempty open $V\subseteq Y$ such that $X_{V}\to V$ has geometrically irreducible fibers.
Lemma 3.10 ([stacks-project],Tag 0559).
Let $X\to Y$ be a morphism of schemes. Assume that
  • $Y$ is irreducible with generic point $\eta$,
  • $\dim X_{\eta}=d$,
  • $X\to Y$ is finite type.
Then there exists a non-empty open $V\subseteq Y$ such that the fibers of $X_{V}\to V$ are $d$-dimensional.
Lemma 3.11.
Let $A=\mathbf{F}_{q}[t]$, and set $K:=\mathbf{F}_{q}(t)$. Let $L/K$ be a finite extension and let $B$ denote the integral closure of $A$ in $L$. Then there are infinitely many pairs $(\mathfrak{p}\in\operatorname{Spec}A,\mathfrak{q}\in\operatorname{Spec}B)$ such that $\mathfrak{q}\cap A=\mathfrak{p}$ and $A/\mathfrak{p}\to B/\mathfrak{q}$ is an isomorphism. In this construction, each $\mathfrak{p}$ occurs in at most one pair.
Proof.
By Lemma 6.4, there is an infinite collection $P$ of prime elements $\pi\in A$ such that for each $\pi$, there is a field embedding $\alpha_{\pi}:L\to K_{\pi},$ where $K_{\pi}$ is the $\pi$-adic completion of $K.$ Let $\pi\in P$. Since $B$ is integral over $A$ and $A^{\wedge\pi}$ is integrally closed in $K_{\pi}$, we have $\alpha_{\pi}(B)\subseteq A^{\wedge\pi}$. Consequently, we consider the ring map $\beta_{\pi}:B\xrightarrow{\alpha_{\pi}}A^{\wedge\pi}\to A/\pi$. Since the image of $\beta_{\pi}$ contains $(A/\pi)^{p^{e}}$ and $A/\pi$ is a finite field, $\beta_{\pi}$ is surjective, hence $B/\textup{ker }\beta_{\pi}\cong A/\pi$. As $\pi\in\textup{ker }\beta_{\pi}$, the prime ideal ker $\textup{ker }b_{\pi}$ lies over $(\pi)$, so the infinite set $\{((\pi),\textup{ker }\beta_{\pi}):\pi\in P\}$ satisfies the desired properties. ∎
Lemma 3.12.
Let $A=\mathbf{F}_{q}[t]$, and set $K:=\mathbf{F}_{q}(t)$. Let $\operatorname{\mathfrak{X}}$ be a finite-type $A$-scheme with generic fiber $X$. There are infinitely many primes $\mathfrak{p}\in\operatorname{Spec}A$ such that the irreducible components of the fiber $\operatorname{\mathfrak{X}}_{\mathfrak{p}}$ are geometrically irreducible and $\dim\operatorname{\mathfrak{X}}_{\mathfrak{p}}=\dim X$.
Proof.
Let $L$ be a finite extension of $K$ such that the irreducible components of $X\times_{\operatorname{Spec}K}\operatorname{Spec}L$ are geometrically irreducible. Let $B$ denote the integral closure of $A$ in $L$ and let $\operatorname{\mathfrak{Y}}:=\operatorname{\mathfrak{X}}\times_{\operatorname{Spec}A}\operatorname{Spec}B$. Consider the map $\operatorname{\mathfrak{Y}}\to\operatorname{Spec}A$. The fiber of $\operatorname{\mathfrak{Y}}$ over $K$ is given by
$\operatorname{\mathfrak{Y}}\times_{\operatorname{Spec}A}K\cong\operatorname{\mathfrak{X}}\times_{\operatorname{Spec}A}\operatorname{Spec}(B\otimes_{A}K)\cong\operatorname{\mathfrak{X}}\times_{\operatorname{Spec}A}\operatorname{Spec}L.$
Consequently, we denote this fiber by $X_{L}$.
Applying Lemma 3.9 once for each irreducible component of $X_{L}$, there is a nonempty open set $U\subseteq\operatorname{Spec}A$ such that the fibers of $\operatorname{\mathfrak{Y}}$ over each $s\in U$ decompose into geometrically irreducible components. Moreover, by Lemma 3.10, there exists a nonempty open set $V\subseteq\operatorname{Spec}A$ such that $\dim\operatorname{\mathfrak{X}}_{s}=\dim X$ for all $s\in V$.
Apply Lemma 3.11 to the extension $K/L$ to obtain an infinite set $Q$ of pairs $\{(\mathfrak{p},\mathfrak{q})\}$. Since $A$ is integral and 1-dimensional, the complement of $U\cap V$ in $\operatorname{Spec}A$ is finite, so we may replace $Q$ by an infinite subset $Q^{\prime}\subseteq Q$ such that $\mathfrak{p}\in U\cap V$ for all $(\mathfrak{p},\mathfrak{q})\in Q^{\prime}$. For each $(\mathfrak{p},\mathfrak{q})\in Q^{\prime}$, factorize $\mathfrak{p}B=\mathfrak{q}^{e_{\mathfrak{p}}}J_{\mathfrak{p}}$ for some ideal $J_{\mathfrak{p}}\subseteq B$ coprime to $\mathfrak{q}$. The reduced fiber of $\operatorname{\mathfrak{Y}}$ over $\operatorname{Spec}A/\mathfrak{p}$ is given by
$(\operatorname{\mathfrak{Y}}_{\mathfrak{p}})_{\text{red}}$ $=(\operatorname{\mathfrak{Y}}\times_{\operatorname{Spec}A}\operatorname{Spec}(A/\mathfrak{p}))_{\text{red}}$
$\cong(\operatorname{\mathfrak{X}}\times_{\operatorname{Spec}A}(\operatorname{Spec}B/\mathfrak{q}^{e_{\mathfrak{p}}}\sqcup\operatorname{Spec}B/J_{\mathfrak{p}})_{\text{red}}$
$\cong(\operatorname{\mathfrak{X}}_{\mathfrak{p}})_{\text{red}}\sqcup(\operatorname{\mathfrak{X}}\times_{\operatorname{Spec}A}\operatorname{Spec}B/J_{\mathfrak{p}}))_{\text{red}}$
As $\mathfrak{p}\in U$, the irreducible components of $\operatorname{\mathfrak{Y}}_{\mathfrak{p}}$ are geometrically irreducible. In particular, as $(\operatorname{\mathfrak{X}}_{\mathfrak{p}})_{\text{red}}$ is a connected component of $(\operatorname{\mathfrak{Y}}_{\mathfrak{p}})_{\text{red}}$, the irreducible components of $(\operatorname{\mathfrak{X}}_{\mathfrak{p}})_{\text{red}}$ are geometrically irreducible, hence the same holds for $\operatorname{\mathfrak{X}}_{\mathfrak{p}}$. Moreover, by assumption that $\mathfrak{p}\in V$, we have $\dim\operatorname{\mathfrak{X}}_{\mathfrak{p}}=\dim X$. ∎
Combining the above lemma with the two point-counting estimates, we have the following.
Lemma 3.13.
Let $A=\mathbf{F}_{q}[t]$, and set $K:=\mathbf{F}_{q}(t)$. Let $\operatorname{\mathfrak{X}}$ be an affine, finite-type $A$-scheme with generic fiber $X$. Then we have
$\limsup_{\mathfrak{p}\in\max\operatorname{Spec}A,|A/\mathfrak{p}|\to\infty}\frac{\log|\operatorname{\mathfrak{X}}_{\mathfrak{p}}(A/\mathfrak{p})|}{\log|A/\mathfrak{p}|}=\dim X.$
Proof.
By Lemma 3.10, we have $\dim\operatorname{\mathfrak{X}}_{\mathfrak{p}}=\dim X$ for all but finitely many primes $\mathfrak{p}$. As the complexity of $\operatorname{\mathfrak{X}}_{\mathfrak{p}}$ is upper-bounded by the complexity of $X$, by Lemma 3.5 we have
$\limsup_{\mathfrak{p}\in\max\operatorname{Spec}A,|A/\mathfrak{p}|\to\infty}\frac{\log|\operatorname{\mathfrak{X}}_{\mathfrak{p}}(A/\mathfrak{p})|}{\log|A/\mathfrak{p}|}\leq\dim X.$
In other direction, Lemma 3.12 implies that for infinitely many primes $\mathfrak{p}\in\operatorname{Spec}A$, $\dim\operatorname{\mathfrak{X}}_{\mathfrak{p}}=\dim X$ and $\operatorname{\mathfrak{X}}_{\mathfrak{p}}$ has a top-dimensional geometrically irreducible component. By Lemma 3.8 we have the following asymptotic, which completes our proof.
$\frac{\log|\operatorname{\mathfrak{X}}_{\mathfrak{p}}(|A/\mathfrak{p}|)|}{\log|A/\mathfrak{p}|}\geq\dim X+\frac{\log(1+O_{X}(|A/\mathfrak{p}|^{-1/2}))}{\log|A/\mathfrak{p}|}.$
The proof of 3.7 is an immediate consequence:
Proof.
Apply Lemma 3.13 to $S_{F}$ to obtain
$\limsup_{|A/\mathfrak{p}|\to\infty}\frac{\log|S_{F\otimes_{A}A/\mathfrak{p}}(A/\mathfrak{p})|}{\log|A/\mathfrak{p}|}=\dim S_{F\otimes_{A}K}.$
Subtract both sides from $n(d-1)$ to conclude the result. ∎
4. Proof of Theorem 1.2 in positive characteristic
Fix a $d$-tensor $F$ over $K=\mathbf{F}_{q}(t)$. To bound the arithmetic rank of $F$ in terms of the geometric rank, one proceeds as follows:
  1. (1)
    First, change bases so that $F$ has coefficients in $A=\mathbf{F}_{q}[t]$. The variety
    $\mathfrak{S}_{F}=(F(x,\cdot)\equiv 0)\subseteq\mathbb{A}_{A}^{n(d-1)}$
    has generic fiber $S_{F}$, giving an integral model for $S_{F}$ over $A$.
  2. (2)
    Using 3.7, one shows that many fibers of $\mathfrak{S}_{F}$ over $\operatorname{Spec}A/\mathfrak{p}$ have many $A/\mathfrak{p}$-rational points.
  3. (3)
    Next, using a scaling result, one can promote some of these $A/\mathfrak{p}$-points of $\mathfrak{S}_{F}$ to $A$-points of small height.
  4. (4)
    Using Noether normalization, these $A$-points of small height give a lower bound on $\dim S_{F}$ in terms of $\textnormal{grk}(F)$, which in turn yields the upper bound $\textnormal{ark}(F)\leq(d-1)\textnormal{grk}(F)$.
  5. (5)
    For finite fields, we extend to $\mathbf{F}_{q}(t)$, apply the above argument, and then bound the size of fibers of the map evaluating at $t=0.$
By induction on the transcendence degree, one can deduce a bound for $\mathbf{F}_{q}(t_{1},\dots,t_{r})$. We begin with a lemma, inspired by a covering argument in [Sch], that will be crucial for step (3).
4.1. Infinite Fields
Lemma 4.1 (Scaling).
Let $H,H^{\prime}$ be finite abelian groups and $G:H^{d-1}\to H^{\prime}$ a multi-linear map. Suppose $\varphi:H\to L$ is a group homomorphism and define
$N^{y}(G)=|\{x\in H^{d-1}:G(x)=0,\varphi(x_{i})=y_{i}\}|$
for any $y\in L^{d-1}.$ Then
(6) $N^{y}(G)\leq N^{0}(G).$
In particular, for any subgroup $H_{0}\subset H,$ we have
|$|\{x\in H^{d-1}:G(x)=0\}|\leq[H:H_{0}]^{d-1}\cdot\left|\{x\in H_{0}^{d-1}:G(x)=0\}\right|.$
Proof.
The statement about subgroups is deduced by taking $\varphi$ to be the projection $\varphi:H\to H/H_{0}$ and then summing over all $y\in(H/H_{0})^{d-1}.$ We prove inequality (6) by induction on $d.$ For the base case $d=2,$ there are two possibilities. If $N^{y}(G)=0$, then there is nothing to prove. Otherwise, there exists some $x_{0}\in H$ with $G(x_{0})=0$ and $\varphi(x_{0})=y.$ In this case,
$N^{y}(G)=\left|x_{0}+\{x\in H:G(x)=0,\varphi(x)=0\}\right|=N^{0}(G).$
For general $d>2,$ writing $y=(y_{1},\dots,y_{d-1})$, we have
$N^{y}(G)$ $=\sum_{x\in H,\ \varphi(x_{d-1})=y_{d-1}}N^{(y_{1},\ldots,y_{d-2})}(G(\cdot,x))\leq\sum_{x\in H,\ \varphi(x_{d-1})=y_{d-1}}N^{0}(G(\cdot,x))$
$=\sum_{x^{\prime}\in H^{d-2},\ \varphi(x^{\prime}_{i})=0}N^{y_{d-1}}(G(x^{\prime},\cdot))\leq\sum_{x^{\prime}\in H^{d-2},\ \varphi(x^{\prime}_{i})=0}N^{0}(G(x^{\prime},\cdot))=N^{0}(G),$
where the first inequality follows from the inductive hypothesis and the second one from the base case $d=2.$
We now define a quantity which refines the notion of arithmetic rank of a $d$-tensor over a rational function field.
Definition 4.2.
Let $A=\mathbf{F}_{q}[t_{1},\dots,t_{r}]$ and let $K=\mathbf{F}_{q}(t_{1},\dots,t_{r})$. Let $F:(A^{n})^{d}\to A$ be a multi-linear form. Define $G:(A^{n})^{d-1}\to\operatorname{Hom}_{A}(A^{n},A)$ by $G(x)=F(x,\cdot)$. For $R_{1},\dots,R_{r}>0$, let
$N_{R_{1},\dots,R_{r}}(G)=|\{x\in(A^{n})^{d-1}:G(x)=0,\ \deg_{t_{i}}(x)<R_{i}\text{ for all }1\leq i\leq r.\}|$
We then define the following invariant of $F$:
$\gamma_{q}(F)=n(d-1)-\limsup_{R_{1},\dots,R_{r}\to\infty}\frac{\log_{q}N_{R_{1},\dots,R_{r}}(G)}{R_{1}\dots R_{r}},$
where the limit is taken only over sequences composed of integers.
Proposition 4.3.
Let $A,K,F$ and $G$ be as in Definition 4.2. Then $\gamma_{q}(F)\leq(d-1)^{r}\textnormal{grk}(F\otimes_{A}K).$
Before proving this lemma, we will show how it implies Theorem 1.2 for rational function fields over $\mathbf{F}_{q}$.
Proof of Theorem 1.2.
Let $K=\mathbf{F}_{q}(t_{1},\dots,t_{r})$ for $r\geq 1$. A standard application of Noether normalization (Lemma 6.1) gives us $\textnormal{ark}(F)\leq\gamma_{q}(F)\leq(d-1)^{r}\textnormal{grk}(F)$. ∎
Now, we return to the proof of 4.3.
Proof.
Write $c=\textnormal{grk}(F\otimes_{A}K).$ For the base case $r=1$, we apply 3.7 to $F$. Let $\mathfrak{p}$ be a prime of $A$ such that $\textnormal{ark}_{A/\mathfrak{p}}(F\otimes_{A}A/\mathfrak{p})\leq c+\varepsilon$. Let $G_{\mathfrak{p}}:(A^{n})^{d-1}\to(A/\mathfrak{p})^{n}$ denote the map obtained by reducing the coefficients of $G$ modulo $\mathfrak{p}$. For any $l>0$, let $H_{l}\subseteq A^{n}$ denote the subgroup of vectors of polynomials of degree strictly less than $l$.
Let $e=\deg\mathfrak{p}$. By lifting each zero $\overline{x}\in((A/\mathfrak{p})^{n})^{d-1}$ of $G\otimes_{A}A/\mathfrak{p}$ to the unique representative $x\in(H_{e})^{d-1}$ of $G_{\mathfrak{p}}$, we get
(7) $N_{e}(G_{\mathfrak{p}})\geq q^{e(n(d-1)-c-\varepsilon)}.$
By Lemma 4.1, for any $0<\sigma<1,$
(8) $N_{e}(G_{\mathfrak{p}})\leq[H_{e}:H_{\sigma e}]^{d-1}N_{{\sigma e}}(G_{\mathfrak{p}})=q^{(e-\lceil\sigma e\rceil)n(d-1)}N_{{\sigma e}}(G_{\mathfrak{p}})$
Combining Equation 7 and Equation 8 yields
$N_{{\sigma e}}(G_{\mathfrak{p}})\geq q^{\lceil\sigma e\rceil(n(d-1)-\frac{c+\varepsilon}{\sigma})}.$
For any vector $x$ with entries in $A,$ let $\lVert x\rVert$ denote the maximum degree of the entries of $x$. There exists a constant $D$ depending on the coefficients of $G$ such that
$\lVert G(x)\rVert\leq D+(d-1)\lVert x\rVert.$
Consequently, when $D+(d-1)\lVert x\rVert<e$, we have $G_{\mathfrak{p}}(x)=0\iff G(x)=0$. It follows that for $\sigma<\frac{1}{d-1}$ and sufficiently large $e,$ we have
(9) $N_{\lceil\sigma e\rceil}(G)=N_{\sigma e}(G)\geq q^{\lceil\sigma e\rceil(n(d-1)-\frac{c+\varepsilon}{\sigma})}.$
Taking $\varepsilon\searrow 0$ and $\sigma\nearrow\frac{1}{d-1}$, we conclude that $\gamma_{q}(F)\leq(d-1)c$ as desired.
Suppose instead $r\geq 2$. By Lemma 3.10, there exists a nonempty open subset $U\subseteq\operatorname{Spec}A$ such that $\textnormal{grk}(F\otimes_{A}\kappa(\mathfrak{p}))=c$ for every $\mathfrak{p}\in U,$ where $\kappa(\mathfrak{p})=\text{Frac}(A/\mathfrak{p})$ is the residue field at $\mathfrak{p}.$ The principal prime ideals in $A$ of the form $(\alpha(t_{r}))$ form a dense subset of $\operatorname{Spec}A$, so $U$ contains infinitely many of these primes. In particular, let $\mathfrak{p}$ be any such prime with $\textnormal{grk}(F\otimes_{A}\kappa(\mathfrak{p}))=c,$ and write $e=\deg\mathfrak{p}$. As in the base case, let $G_{\mathfrak{p}}:(A^{n})^{d-1}\to(A/\mathfrak{p})^{n}$ denote the map obtained by reducing the coefficients of $G$ modulo $\mathfrak{p}$.
By the induction hypothesis, we have $\gamma_{q^{e}}(F\otimes_{A}A/\mathfrak{p})\leq(d-1)^{r-1}c.$ We may therefore choose $R_{1},\dots,R_{r-1}$ arbitrarily large with
(10) $N_{R_{1},\dots,R_{r-1},e}(G_{\mathfrak{p}})\geq q^{eR_{1},\dots,R_{r-1}(n(d-1)-(d-1)^{r-1}c-\varepsilon)}.$
For $l_{1},\ldots,l_{r}>0,$ let $H_{l_{1},\ldots,l_{r}}\subset A^{n}$ be the additive subgroup of vectors with entries whose degree in $t_{i}$ is less than $l_{i}$ for all $i.$ Applying Lemma 4.1 to the function $G_{\mathfrak{p}}:H_{e}\to(A/\mathfrak{p})^{n}$ with $0<\sigma<\frac{1}{d-1}$ and arguing similarly to the base case, we obtain
(11) $N_{R_{1},\dots,R_{r-1},\lceil\sigma e\rceil}(G)\geq q^{\lceil\sigma e\rceil R_{1},\dots,R_{r-1}\left(n(d-1)-\frac{(d-1)^{r-1}c+\varepsilon}{\sigma}\right)}.$
Taking $\varepsilon\searrow 0$ and $\sigma\nearrow\frac{1}{d-1}$ yields $\gamma_{q}(F)\leq(d-1)^{r}c,$ as desired. ∎
4.2. Finite fields
Lemma 4.4 (Fibers of evaluation at zero).
Let $F:(\mathbf{F}_{q}^{n})^{d}\to\mathbf{F}_{q}$ be a multi-linear map and let $G:(\mathbf{F}_{q}[t]^{n})^{d-1}\to\mathbf{F}_{q}[t]^{n}$ be the multi-linear map given by $G(x)_{i}=F_{\mathbf{F}_{q}(t)}(x,e_{i}).$ Then for any integer $R>1$ we have
$N_{R}(G)\leq|S_{F}(\mathbf{F}_{q})|\cdot N_{R-1}(G),$
where $S_{F}=(F(x,\cdot)\equiv 0).$
Proof.
First note that $x(0)\in S_{F}(\mathbf{F}_{q})$ for any $x\in(\mathbf{F}_{q}[t]^{n})^{d-1}$ with $G(x)=0.$ For any $a\in S_{F}(\mathbf{F}_{q}),$ let
$N_{R}^{a}=|\{x\in S_{F}(\mathbf{F}_{q}[t]),\ \deg(x)<R,x(0)=a\}|.$
By lemma 4.1, we have $N_{R}^{a}(G)\leq N_{R}^{0}(G)=N_{R-1}(G).$ The equality follows from the bijection given by $x\mapsto x/t.$ Summing over $a\in S_{F}(\mathbf{F}_{q})$ yields
$N_{R}(G)=\sum_{a\in S_{F}(\mathbf{F}_{q})}N_{R}^{a}(G)\leq\sum_{a\in S_{F}(\mathbf{F}_{q})}N_{R-1}(G)=|S_{F}(\mathbf{F}_{q})|\cdot N_{R-1}(G).$
Proof of theorem 1.2 for finite fields.
This follows from the fact that
$\textnormal{ark}(F)\leq\gamma_{q}(F_{\mathbf{F}_{q}(t)})\leq(d-1)\textnormal{grk}(F).$
We prove this now. By Lemma 4.4, for any $R\geq 1$, we have
(12) $N_{R}(G)=|N_{1}(G)|\cdot\prod_{j=1}^{R-1}\frac{N_{j+1}(G)}{N_{j}(G)}\leq|S_{F}(\mathbf{F}_{q})|^{R}.$
By 4.3, we have $\gamma_{q}(F)\leq(d-1)\textnormal{grk}(F)$, so for any $\varepsilon>0$, there exists $R\in\mathbf{Z}^{+}$ such that
(13) $N_{R}(G)\geq q^{R(n(d-1)-(d-1)\textnormal{grk}(F)-\varepsilon)}.$
Combining Equation 13 and Equation 12 we obtain
$\textnormal{ark}(F)=\log_{q}(|S_{F}(\mathbf{F}_{q})|)\leq(d-1)\textnormal{grk}(F)+\varepsilon$
for all $\varepsilon>0$, proving the claim.
5. Proof of Theorem 1.2 in characteristic zero
In this section, $F$ denotes a $d$-tensor over $K=\mathbf{Q}(t_{1},\dots,t_{s})$ for some $s\geq 0$. Using the results of the previous section, we will derive the bound $\textnormal{ark}(F)\leq(d-1)^{s+1}\textnormal{grk}(F)$.
  1. (1)
    First, change bases so that $F$ has coefficients in $A=\mathbf{Z}[t_{1},\dots,t_{s}]$. The variety
    $\mathfrak{S}_{F}=(F(x,\cdot)\equiv 0)\subseteq\mathbb{A}_{A}^{n(d-1)}$
    has generic fiber $S_{F}$, giving an integral model for $S_{F}$ over $A$.
  2. (2)
    For $s=0$, the argument is essentially the same as the case of $\mathbf{F}_{q}[t]$ in the previous section, so we assume $s\geq 1$ for the remainder of this outline.
  3. (3)
    For all but finitely many primes $p$, the form $F\otimes_{A}A/p$ has the same geometric rank as $F$. By 4.3, the variety $S_{F\otimes_{A}A/p}$ has many rational points of small height. Similarly to the previous section, many of these points lift to $A$-points of $\mathfrak{S}_{F}.$
  4. (4)
    By Noether normalization, the abundance of $A$-points proves the desired upper bound $\textnormal{ark}(F)\leq(d-1)^{s+1}\textnormal{grk}(F)$.
We begin with a semigroup analogue of lemma 4.1.
Lemma 5.1 (Scaling).
Let $B$ be an abelian group and let $G:(\mathbf{Z}^{n})^{d-1}\to B$ be a multi-linear map. For a positive integer $R$ set
[…]
$Z_{R}(G)$ =|$=|\{x\in(\mathbf{Z}^{n})^{d-1}:G(x)=0,\ -R<x_{j}<R\ \forall j\in[n(d-1)]\}|.$
Then for any positive integer $L$ we have
$N_{LR}(G)\leq L^{n(d-1)}Z_{R}(G).$
Proof.
By induction on $d.$
We begin with the base case $d=2$. First note that for any $b\in B$ the number of integer solutions to $G(x)=b$ in $[0,R)^{n}$ is bounded above by $Z_{R}(G).$ This follows by subtracting a single solution from each of the others. Therefore
$N_{LR}(G)$ $=\sum_{0\leq x_{i}<LR}1_{G(x)=0}=\sum_{\omega\in\{0,1,\ldots,L-1\}^{n}}\sum_{0\leq x_{i}<R}1_{G(x+R\omega)=0}$
$=\sum_{\omega\in\{0,1,\ldots,L-1\}^{n}}\sum_{0\leq x_{i}<R}1_{G(x)=-G(R\omega)}\leq\sum_{\omega\in\{0,1,\ldots,L-1\}^{n}}Z_{R}(G)=L^{n}Z_{R}(G).$
For general $d>2,$ we have
$N_{LR}(G)$ $=\sum_{x\in\mathbf{Z}^{n},0\leq x_{i}<LR}N_{LR}(G(x,\cdot))\leq\sum_{x\in\mathbf{Z}^{n},0\leq x_{i}<LR}L^{n(d-2)}Z_{R}(G(x,\cdot))$
$=L^{n(d-2)}\sum_{y\in\mathbf{Z}^{n(d-2)},-R<y_{i}<R}N_{LR}(G(\cdot,y))$
$\leq L^{n(d-1)}\sum_{y\in\mathbf{Z}^{n(d-2)},-R<y_{i}<R}Z_{R}(G(\cdot,y))=L^{n(d-1)}Z_{R}(G),$
where the first inequality follows from the inductive hypothesis and the second one from the base case $d=2.$
Definition 5.2.
Let $A=\mathbf{Z}[t_{1},\dots,t_{s}]$ and let $K=\mathbf{Q}(t_{1},\dots,t_{s})$. Let $F:(A^{n})^{d}\to A$ be a $d$-tensor. Define $G:(A^{n})^{d-1}\to\operatorname{Hom}_{A}(A^{n},A)$ by $G(x)=F(x,\cdot)$. Let $A_{\geq 0}$ be the additive sub-semigroup of $A$ in which the coefficients of all monomials are non-negative. For $x\in(A^{n})^{d-1}$, let $\lVert x\rVert\in\mathbf{Z}^{\geq 0}$ denote the largest absolute value of any integer coefficient appearing in any term of $x$. For $L,R_{1},\dots,R_{s}\in\mathbf{Z}^{+}$, we define
$N_{L,R_{1},\dots,R_{s}}(G)=|\{x\in(A_{\geq 0}^{n})^{d-1}:G(x)=0,\ \lVert x\rVert<L,\ \deg_{t_{i}}(x)<R_{i}\text{ for all }1\leq i\leq s\}|$
and similarly define
$Z_{L,R_{1},\dots,R_{s}}(G)=|\{x\in(A^{n})^{d-1}:G(x)=0,\ \lVert x\rVert<L,\ \deg_{t_{i}}(x)<R_{i}\text{ for all }1\leq i\leq s\}|.$
We then define the following invariants of $F$:
$\gamma_{0}(F)=n(d-1)-\limsup_{L,R_{1},\dots,R_{s}\to\infty}\frac{\log_{L}Z_{L,R_{1},\dots,R_{s}}(G)}{R_{1}\dots R_{s}}.$
[…]
$\delta_{0}(F)=n(d-1)-\limsup_{L,R_{1},\dots,R_{s}\to\infty}\frac{\log_{L}N_{L,R_{1},\dots,R_{s}}(G_{L})}{R_{1}\dots R_{s}}.$
Note that this definition makes sense even if $s=0$, as the empty product equals 1.
Lemma 5.3.
Let $A,K,F,G$ be as in Definition 5.2. Then $\gamma_{0}(F)\leq(d-1)\delta_{0}(F)$.
Assuming this lemma, we are now able to prove Theorem 1.2 for fields of characteristic zero.
Proof.
A standard application of Noether normalization (Lemma 6.2) shows that $\textnormal{ark}(F)\leq\gamma_{0}(F)$. By Lemma 5.3, it suffices to prove that $\delta_{0}(F)\leq(d-1)^{s}\textnormal{grk}(F)$. For $s=0$, this follows from [KMZ20], Theorem 8.1. For $s\geq 1$, we note by Lemma 3.10 that for all but finitely many primes $p\in\mathbf{Z}$, we have $\textnormal{grk}(F\otimes_{A}\kappa(p))=\textnormal{grk}(A)$. Consequently, the bound $\delta_{0}(F)\leq(d-1)^{s}\textnormal{grk}(F)$ follows from 4.3 for all $s\geq 1$. ∎
To conclude this section, we now prove Lemma 5.3.
Proof.
Choose $L,R_{1},\dots,R_{s}\in(\mathbf{Z}^{+})^{s+1}$ such that
(14) $N_{L,R_{1},\dots,R_{s}}(G_{L})\geq L^{R_{1}\dots R_{s}(n(d-1)-\delta_{0}(F)-\varepsilon)}.$
Let $R=R_{1}\dots R_{s}$. The basis given by monomials yields an isomorphism $\mathbf{Z}[t_{1},\dots,t_{s}]_{\deg t_{i}<R_{i}}\cong\mathbf{Z}^{R}$ of abelian groups, which sends $G_{L}$ to a multi-linear map $\tilde{G}_{L}:(\mathbf{Z}^{Rn})^{d-1}\to(A/L)^{n}.$ Under this identification, $\lVert x\rVert$ denotes the largest absolute value of any entry of the corresponding vector in $(\mathbf{Z}^{Rn})^{d-1}.$ For $10<\sigma<$ 1, Equation 14 together with Lemma 5.1 gives us
$L^{R(n(d-1)-\delta_{0}(F)-\varepsilon)}\leq N_{L}(\tilde{G}_{L})\leq\lceil L^{1-\sigma}\rceil^{Rn(d-1)}Z_{\lceil L^{\sigma}\rceil}(\tilde{G}_{L})\ll_{d,n,R}L^{R(1-\sigma)n(d-1)}Z_{\lceil L^{\sigma}\rceil}(G_{L}).$
In particular, we have
(15) $Z_{\lceil L^{\sigma}\rceil}(G_{L})\gg_{d,n,R}L^{R\sigma\left(n(d-1)-\frac{\delta_{0}(F)+\varepsilon}{\sigma}\right)}\gg_{d,n,R}\lceil L^{\sigma}\rceil^{Rn(d-1)-\frac{\delta_{0}(F)+\varepsilon_{\vec{v}}}{\sigma}}.$
For any $x\in(\mathbf{Z}[t_{1},\dots,t_{s}]_{\deg t_{i}<R_{i}}^{n})^{d-1}$ satisfying $\|x\|<\lceil L^{\sigma}\rceil$, we have $\lVert G(x)\rVert\ll_{F}L^{\sigma(d-1)}$. Therefore, for $L$ sufficiently large, $\sigma<\frac{1}{d-1}$ and $\|x\|<\lceil L^{\sigma}\rceil$, we have
$G_{L}(x)=0\mod L\iff G(x)=0.$
Consequently, Equation 15 yields
$\gamma_{0}(F)\leq\frac{\delta_{0}(F)+\varepsilon}{\sigma}.$
Taking the limit $\varepsilon\searrow 0$ and $\sigma\nearrow\frac{1}{d-1}$ completes the proof. ∎
6. Appendix
The results contained within this appendix are standard and similar to existing results. Nevertheless, we detail their proofs here for the sake of self-containedness and completeness.
6.1. A Consequence of Noether Normalization
Lemma 6.1 (Height point growth in positive characteristic).
Let $A=\mathbf{F}_{q}[t_{1},\dots,t_{r}],K=\mathbf{F}_{q}(t_{1},\dots,t_{r})$, and let $X\subset\mathbb{A}^{n}$ be an affine variety defined over $K.$ We define a height function on points of $A^{n}$ by $H(x)=(\deg_{t_{1}}x,\ldots,\deg_{t_{r}}x).$ For $(R_{1},\dots,R_{r})\in\mathbb{N}^{r}$, set
$X(R_{1},\dots,R_{r})=\{x\in X(A):H(x)_{i}<R_{i}\text{ for all }1\leq i\leq r\}.$
Then
$\dim(X)\geq\limsup_{R_{1},\dots,R_{r}\to\infty}\frac{\log_{q}|X(R_{1},\dots,R_{r})|}{R_{1}\dots R_{r}}.$
Proof.
The key is that we have
$|\mathbb{A}^{b}(R_{1},\dots,R_{r})|=|\mathbb{A}^{1}(R_{1},\dots,R_{r})|^{b}=q^{bR_{1}\dots R_{r}}.$
By Noether normalization, there exists a finite map $\pi:X\to\mathbb{A}^{b}$ for $b=\dim X$. Since $A$ is infinite, this map can be taken linear with coefficients in $A$. Consequently, there exists $C>0$ such that
$\pi(X(R_{1},\dots,R_{s}))\subseteq\mathbb{A}^{b}(R_{1}+C,\dots,R_{r}+C).$
Taking cardinalities of both sides, we obtain
$|X(R_{1},\dots,R_{r})|\leq\deg(\pi)q^{b(R_{1}+C)\dots(R_{r}+C)}.$
Taking $\log_{q}$, dividing by $R_{1}\ldots R_{r}$ and then taking the limit $R_{1},\dots,R_{r}\to\infty$ completes the proof. ∎
Lemma 6.2 (Height point growth in characteristic zero).
Let $A=\mathbf{Z}[t_{1},\dots,t_{s}],K=\mathbf{Q}(t_{1},\ldots,t_{s})$, and let $X\subset\mathbb{A}^{n}$ be an affine variety defined over $K.$ We’ll define a height function on points of $A^{n}$ by $H(x)=(L,\deg_{t_{1}}x,\dots,\deg_{t_{s}}x)$, where $L$ denotes the maximum of the absolute values of the coefficients of the $x.$ For $(L,R_{1},\dots,R_{s})\in\mathbb{N}^{s+1}$, set
$X(L,R_{1},\dots,R_{s})=\{x\in X(A):H(x)_{0}<L,H(x)_{i}<R_{i}\text{ for all }1\leq i\leq s\}.$
Then
$\dim(X)\geq\limsup_{L,R_{1},\dots,R_{s}\to\infty}\frac{\log_{L}(X(L,R_{1},\dots,R_{s}))}{R_{1}\dots R_{s}}\geq a.$
Proof.
The key is that
$|\mathbb{A}^{b}(L,R_{1},\dots,R_{s})|=|\mathbb{A}^{1}(L,R_{1},\dots,R_{s})|^{b}=(2L-1)^{bR_{1}\dots R_{s}}\leq(2L)^{bR_{1}\dots R_{s}}.$
By Noether normalization, there exists a finite map $\pi:X\to\mathbb{A}^{b}$ for $b=\dim X$. Since $A$ is infinite, this map can be taken linear with coefficients in $A$. Consequently, there exists $C>0$ such that
$\pi(X(L,R_{1},\dots,R_{s}))\subseteq\mathbb{A}^{b}(CL,R_{1}+C,\dots,R_{s}+C).$
Taking cardinalities of both sides, we obtain
$|X(L,R_{1},\dots,R_{s})|\leq\deg(\pi)(2CL)^{b(R_{1}+C)\dots(R_{s}+C)}.$
Taking $\log_{L}$, dividing by $R_{1}\dots R_{s}$ and taking the limit $L,R_{1},\dots,R_{s}\to\infty$ completes the proof. ∎
6.2. Embeddings into Local Fields
In [C76], Cassels constructs embeddings from a finitely generated extension of $\mathbf{Q}$ into $\mathbf{Q}_{p}$ for infinitely many primes $p$. For our purposes, we require a weaker result for finite extensions of $\mathbf{F}_{q}(t)$. These arguments are minor modifications of Cassels’s own.
Lemma 6.3.
Let $A=\mathbf{F}_{q}[t]$. Let $G(X)\in A[X]$ be a non-constant function. Then $G$ has a solution mod $\pi$ for infinitely many prime elements $\pi\in A$.
Proof.
Cassels gives an elementary proof of this result in [C76A] for $A=\mathbb{Z}$. The proof for $\mathbf{F}_{q}[t]$ is identical, with only one claim requiring special justification in this case: if $G\in\mathbf{F}_{q}[t][X]$ such that $G(\mathbf{F}_{q}[t])\subseteq\mathbf{F}_{q}[t]^{\times}=\mathbf{F}_{q}^{\times}$ then $G\in\mathbf{F}_{q}[t]$. To see this, note that if $G=G_{s}X^{s}+\dots+G_{0}$ and $s>0$, then $G(t^{n})$ is non-constant for any $n$ larger than $\max_{0\leq i\leq s-1}\deg(G_{i})$. ∎
Lemma 6.4.
Let $A=\mathbf{F}_{q}[t],\ K=\mathbf{F}_{q}(t)$ and let $L/K$ be a finite extension. Let $e\geq 0$ such that $L^{p^{e}}\subseteq K^{\text{sep}}$. Then there are infinitely many primes $\pi\in A$ which admit a field embedding $\alpha_{\pi}:L\to K_{\pi}$ extending the Frobenius iterate $F^{e}:A\to A$.
Proof.
Start with the case that $L/K$ is separable and write $L=K(\beta)$ using the primitive element theorem. Suppose $\beta$ satisfies minimal equation $H(z)=H_{s}z^{s}+\dots+H_{0}$ over $K$. Clear denominators so that $H_{i}\in A$. Let $P^{\prime}$ denote the set of prime elements $\pi\in A$ such that $H(z)=0$ has a solution mod $\pi$. Remove from $P^{\prime}$ all primes which divide $(H_{s}\Delta(H))$. By Lemma 6.3, the resulting set $P$ is still infinite. Let $\pi\in P$. By the fact that
$\Delta H\not\equiv 0\mod\pi,$
it follows that $H$ has a simple root mod $\pi$. By Hensel’s lemma, it follows that $H$ has a root $\eta\in A^{\wedge\pi}$. We then define $\alpha_{\pi}(\beta)=\eta$, which gives a well-defined embedding $\alpha_{\pi}:L\to K_{\pi}$ extending the identity map $A\to A$. In the case that $L/K$ is inseparable, we apply the above argument to $L^{p^{e}}/K$ for some $e$ such that $L^{p^{e}}\subseteq K^{\text{sep}}$ and precompose with $F^{e}:L\to L^{p^{e}}$. ∎
[…]